# T2VEval: Benchmark Dataset and Objective Evaluation Method for T2V-generated Videos

### v5

@@ Abstract
Index Terms:
Text-to-Video, Quality Assessment, Evaluation benchmark, Objective Evaluation
I INTRODUCTION
Text-to-video (T2V) technology has made significant advancements in the last two years and attracted growing attention from both the research community and the general public. As an important branch of Generative Artificial Intelligence (GenAI), T2V models [blattmann2023align, blattmann2023stable, ho2022imagen, singer2022make, wu2021godiva, wu2022nuwa, hong2022cogvideo, villegas2022phenaki, zhou2022magicvideo, khachatryan2023text2video, luo2023videofusion, he2022latent, wang2023modelscope, yang2024cogvideox, zhang2024show, wang2024recipe, videoworldsimulators2024] have developed rapidly recently. Contemporary T2V models, such as Sora, Runway Gen-3, PixVerse, Kling, etc. have enabled video generation from free-form textual prompts and demonstrated impressive visual fidelity and alignment with the prompts. However, as illustrated in Figure 1, the T2V-generated videos still exhibit several notable deficiencies that cannot be overlooked. These shortcomings can significantly detract from the user’s viewing experience. Therefore, developing a comprehensive and effective text-to-video quality assessment strategy is essential to ensure an optimal user experience and to guide the optimization of T2V models.
In recent years, numerous studies have focused on establishing T2V evaluation benchmarks and automated metrics [liu2024fetv, liu2024evalcrafter, wu2024towards, huang2024vbench, zhang2024rethinking, kou2024subjective, sun2024t2v] for assessing the text-to-video quality across dimensions such as visual quality, temporal consistency, and text-video alignment. Although these approach offers a more detailed representation of the generative abilities of T2V models, it may lead to metrics that are less applicable and generalizable due to their over-specialization to specific evaluation dimensions. For instance, readability assessments of the text-related elements in T2V-generated videos become redundant in contexts where textual content is absence or not the primary focus. Furthermore, many works design automated metrics by directly employing models such as DOVER[wu2023exploring] and RAFT[teed2020raft], which are trained and tested on datasets consisting of human-made videos rather than AI-generated videos. This operation ignores the inherent statistical distribution differences between these two classes of videos. This practice overlooking the inherent statistical distribution differences between real videos and text-generated videos. In addition, some studies, similar to traditional video and image quality assessments, use a single score to represent the quality of T2V-generated videos. Although this approach simplifies the evaluation process and enhances metric applicability, it can easily overlook the strengths and weaknesses of certain T2V models in specific generation tasks.
Figure 1: Examples of defects in T2V-generated videos. (a) As shown in the yellow matrix part of the picture, there is obvious distortion similar to ”block effect” in the picture. (b) As shown in the red matrix part of the picture, the camel’s limbs and head are ”broken”. (c) The prompt describes a cheetah chasing an antelope. However, the generated video only has a cheetah, no antelope, and the state of ”chasing” is not simulated.
@@ I INTRODUCTION
Based on the T2VEval-Bench dataset, we further designed T2VEval, an objective evaluation method for T2V-generated. Previous works, such as FETV [liu2024fetv], EvalCrafter [liu2024evalcrafter], and VBench [huang2024vbench], have developed dedicated evaluation metrics for each evaluation dimension. However, these metrics generally lack broad applicability, and they directly employ models trained on datasets with human-made videos, thereby overlooking the inherent distributional differences between human-made videos and AI-generated videos. Other studies like T2VQA[kou2024subjective] and AIGC-VQA [lu2024aigc] have implemented end-to-end evaluations of videos, but they only consider video quality and text-video alignment. Therefore, we designed a multi-branch fusion evaluation scheme, which includes a technical quality encoding branch, a video realness encoding branch, and a text-video alignment encoding branch. The technical quality encoder is implemented based on the Swin-3D network [liu2022video], the video realness encoder utilizes the ConvLexNet-3D network [liu2022convnet], and the text-video alignment is achieved using the BLIP network [li2022blip]. Considering that technical quality and realness rely solely on visual perception, we prioritize the fusion of quality and realness features before integrating them with text-video alignment features. Additionally, we adopted a progressive training strategy that ensures the independent effectiveness of each branch while promoting collaboration among the branches. Experimental results show that T2VEval effectively assesses the performance of different T2V-generated videos and has broad potential applications, achieved state-of-the-art performance on both T2VEval-Bench and T2VQA-DB[kou2024subjective].
II RELATED WORK
II-A Subjective evaluation and dataset for T2V Generation
@@
Chivileva et al. [chivileva2023measuring] constructed the first T2V generated video evaluation dataset. They obtained 201 textual prompts by combining ChatGPT and manual writing, covering topics such as famous figures, regions, and special festivals. Five T2V models, namely Aphantasia, Text2Video-Zero, T2VSynthesis, Tune-a-Video, and Video Fusion, were selected to generate 1005 videos. Subsequently, 24 subjects were invited to subjectively score the generated videos from the perspectives of naturalness and consistency. Liu et al. proposed FETV [liu2024fetv]. They designed a set of diverse and fine-grained textual prompts and classified them into three aspects: main content, control attributes, and text complexity. At the same time, a time category was specifically designed for text-to-video, covering spatial and temporal dimensions. On this basis, the authors selected four T2V models, including CogVideo, Text2Video-zero, ModelScopeT2V, and ZeroScope, and conducted subjective evaluations in terms of static quality, temporal quality, overall alignment, and fine-grained alignment. EvalCrafter [liu2024evalcrafter] first constructed a comprehensive textual prompt corpus covering daily objects, attributes, and actions. It specified the metadata of common elements in the real world and expanded them using large language models like ChatGPT to obtain complete text prompts. It also enriched and expanded the prompts through those written by real-world users and from text-to-image (T2I) datasets. Subsequently, 2500 videos were generated on 5 T2V models, and subjective evaluations were carried out in terms of video quality, text-video consistency, motion quality, and time consistency. Wu et al. [wu2024towards] constructed the TVGE dataset. They adopted the prompt set proposed in EvalCrafter [liu2024evalcrafter] and generated 2543 videos on 5 T2V models. Ten professional scorers were invited to score the videos in terms of video quality and video-text consistency. Huang et al. [huang2024vbench] believed that the existing evaluation benchmarks could not accurately reflect the highlights and problems of different generation models. Based on this, they proposed Vbench, which further refined the evaluation dimensions of T2V. Sixteen fine-grained evaluation dimensions for generated videos were designed, including subject identity inconsistency, motion smoothness, time flicker, and spatial relationship. One hundred text prompts were designed for each evaluation dimension. In addition, the authors also evaluated T2V models for different content categories, such as people, animals, and landscapes. A total of 8 content categories were considered, and 100 prompts were designed for each category. Based on this, the authors conducted subjective evaluations on 4 T2V models and collected corresponding 16-dimensional scores. Li et al. [zhang2024benchmarking] constructed the LGVQ dataset, which contains 468 text prompts covering combinations of different foreground contents, background contents, and motion states. Six T2V models were used to generate 2808 videos, and 60 subjects were invited to perceptually score the spatial quality, temporal quality, and text-video alignment of each video. Based on the LGVQ dataset, the authors carried out benchmark tests using 14 IQA methods, 16 VQA methods, and 9 methods based on CLIP and visual question answering. Zhang et al. [zhang2024rethinking] pointed out that the existing subjective evaluation benchmarks for T2V generally have limitations in reliability, reproducibility, and practicability. In view of this, this paper proposes the T2VHE benchmark, which specifies the evaluation dimensions (including video quality, temporal quality, motion quality, text alignment, etc.) used in T2V evaluation, the annotator training process and example descriptions, and a user-friendly GUI interface. The authors further proposed a dynamic scoring scheme. By combining the automatic scoring method to pre-annotate the videos, the amount of manually annotated data was subsequently reduced by 50%. Kou et al. [kou2024subjective] believed that the current T2V evaluation benchmarks have overly complex divisions of evaluation dimensions and lack a simple and effective evaluation index. Therefore, the authors constructed T2VQA-DB and invited 24 subjects to score 10,000 text-generated videos. The subjects were required to give an overall score considering video quality and text-video consistency.
@@
All of the above works focus on constructing general subjective evaluation benchmarks for text-to-video generation. At the same time, there are also works that conduct subjective evaluations for specific video content types or evaluation angles. Chen et al. [chen2024gaia] constructed the GAIA dataset for action quality assessment of generated videos from the perspective of actions in the generated videos, covering various full-body, hand, and facial actions. TC-Bench [feng2024tc] focuses on evaluating the temporal composition and consistency of each element in the generated videos. T2VBench [ji2024t2vbench] and DEVIL [liao2024evaluation] evaluate from the perspective of the dynamic degree of text-to-video. T2V-CompBench [sun2024t2v] evaluates T2V models from seven dimensions such as dynamic attribute binding, spatial relationship, and generation computing ability in terms of the ability of T2V models to combine different objects, attributes, and actions. In addition, Miao et al. [miao2024t2vsafetybench] paid special attention to the security risks of text-to-video. They pointed out that the generated videos may contain illegal or unethical content, and previous work lacked a quantitative understanding of security. Therefore, the authors introduced T2VSafetyBench to evaluate T2V models in 12 aspects such as violence, gore, discrimination, political sensitivity, copyright, and trademark infringement.
@@
II-B Objective Metrics for T2V Generation
Chivileva et al. [chivileva2023measuring] proposed the first T2V evaluation index, which assesses the naturalness of the video and the similarity between the text and the video. In the text similarity evaluation, corresponding captions are generated for each video based on the BLIP model. Then, the similarity between the generated captions and the text prompts is calculated using the text encoder of BLIP. For the evaluation of naturalness, a series of handcrafted features are introduced, including texture, sharpness, color, spectrum, entropy, contrast, keypoint detection, and the number and size of speckles. FETV [liu2024fetv] aimed at the problem that previous text-to-video evaluation metrics such as FVD, CLIPSim, and IS could not accurately reflect human subjective perception. It improved FVD and CLIPSim and proposed the index FVD-UMT for evaluating video fidelity and the index UMTScore for evaluating text-video alignment. Wu et al. [wu2024towards] proposed T2VScore to automatically evaluate text-to-video in terms of video quality and text-video consistency. T2VScore consists of two components, T2VScore-A and T2VScore-Q, which evaluate text-video consistency and video quality respectively. Specifically, T2VScore-A uses LLMs to generate question-answer pairs for the prompt and then uses MLLMs to ask questions about the video and answer them. The consistency score is obtained by calculating the answer accuracy. The structure of T2VScore-Q adopts a mixture of experts structure, and the basic architecture adopts Fast-VQA. EvalCrafter [liu2024evalcrafter] believes that traditional metrics such as FVD and IS cannot well fit human perception, and simple metrics cannot achieve effective evaluation. The article divides video quality, consistency, motion quality, and time consistency into fine-grained sub-dimensions, obtaining a total of 17 evaluation dimensions, and designs a dedicated evaluation algorithm for each dimension (for example, to evaluate the ability of the current T2V model to generate text, EvalCrafter directly uses the PaddleOCR toolkit [ma2019paddlepaddle] to extract text from the generated video. Then, the word error rate, normalized edit distance, and character error rate are calculated and averaged). VBench [huang2024vbench] also makes a finer division based on the evaluation of video quality and text-video consistency, finally obtaining 16 evaluation dimensions and corresponding 16 evaluation algorithms. Kou et al. [kou2024subjective] designed the first end-to-end text-to-video evaluation model T2VQA based on the proposed T2VQA-DB dataset. It includes a video fidelity encoding module based on Siwn-3D and a text-video consistency representation module based on BLIP. Finally, the quality score is regressed through a large language model. Lu et al. [lu2024aigc] improved on T2VQA by introducing a progressive training strategy, enabling different branches of the model to gradually learn relevant knowledge and achieving SOTA performance on T2VQA-DB. Li et al. [zhang2024benchmarking] found that the current metrics could not comprehensively and effectively evaluate text-to-video based on the LGVQ benchmark test. Therefore, the authors proposed the UGVQ model, which systematically extracts spatial, temporal, and text-video alignment features by combining the ViT model, SlowFast network, and CLIP model, and obtains multi-dimensional scores through MLP regression.
III T2VEval-Bench: a Multi-Dimensional Benchmark Dataset
III-A Division of Subjective Evaluation Dimensions
Traditional video quality assessment task [kou2024subjective, wang2019youtube, sinno2018large, konvid1k, hosu2017konstanz] typically represents video quality using a single score, this approach may obscure the strengths and weaknesses of T2V models in specific tasks. It fails to provide a detailed analysis of the dimensions where the model performs well or poorly. Existing evaluation approaches targeting T2V models also suffer from limited applicability. To address these challenges, we proposed a multi-dimensional subjective evaluation method. Specifically, we decomposed the evaluation of text-to-video generation into four dimensions:
@@ III-A Division of Subjective Evaluation Dimensions
III-B Compilation and Screening of Prompts

### v6

@@ Abstract
keywords:
[…]
1 Introduction
Text-to-video (T2V) technology has made significant advancements in the last two years and attracted growing attention from both the research community and the general public. As an important branch of Generative Artificial Intelligence (GenAI), T2V models Blattmann et al. (2023b, a); Ho et al. (2022); Singer et al. (2022); Wu et al. (2021, 2022a); Hong et al. (2022); Villegas et al. (2022); Zhou et al. (2022); Khachatryan et al. (2023); Luo et al. (2023); He et al. (2022); Wang et al. (2023); Yang et al. (2024); Zhang et al. (2024a); Wang et al. (2024); Brooks et al. (2024) have developed rapidly recently. Contemporary T2V models, such as Sora, Runway Gen-3, PixVerse, Kling, etc. have enabled video generation from free-form textual prompts and demonstrated impressive visual fidelity and alignment with the prompts. However, as illustrated in Figure 1, the T2V-generated videos still exhibit several notable deficiencies that cannot be overlooked. These shortcomings can significantly detract from the user’s viewing experience. Therefore, developing a comprehensive and effective text-to-video quality assessment strategy is essential to ensure an optimal user experience and to guide the optimization of T2V models.
Figure 1: Examples of defects in T2V-generated videos. (a) As shown in the yellow matrix part of the picture, there is obvious distortion similar to "block effect" in the picture. (b) As shown in the red matrix part of the picture, the camel’s limbs and head are "broken". (c) The prompt describes a cheetah chasing an antelope. However, the generated video only has a cheetah, no antelope, and the state of "chasing" is not simulated.
@@ 1 Introduction
In recent years, numerous studies have focused on establishing T2V evaluation benchmarks and automated metrics Liu et al. (2024b, a); Wu et al. (2024b); Huang et al. (2024); Zhang et al. (2024b); Kou et al. (2024); Sun et al. (2024) for assessing the text-to-video quality across dimensions such as visual quality, temporal consistency, and text-video alignment. Although these approach offers a more detailed representation of the generative abilities of T2V models, it may lead to metrics that are less applicable and generalizable due to their over-specialization to specific evaluation dimensions. For instance, readability assessments of the text-related elements in T2V-generated videos become redundant in contexts where textual content is absence or not the primary focus. Furthermore, many works design automated metrics by directly employing models such as DOVERWu et al. (2023a) and RAFTTeed and Deng (2020), which are trained and tested on datasets consisting of human-made videos rather than AI-generated videos. This operation ignores the inherent statistical distribution differences between these two classes of videos. This practice overlooking the inherent statistical distribution differences between real videos and text-generated videos. In addition, some studies, similar to traditional video and image quality assessments, use a single score to represent the quality of T2V-generated videos. Although this approach simplifies the evaluation process and enhances metric applicability, it can easily overlook the strengths and weaknesses of certain T2V models in specific generation tasks.
@@
Based on the T2VEval-Bench dataset, we further designed T2VEval, an objective evaluation method for T2V-generated. Previous works, such as FETV Liu et al. (2024b), EvalCrafter Liu et al. (2024a), and VBench Huang et al. (2024), have developed dedicated evaluation metrics for each evaluation dimension. However, these metrics generally lack broad applicability, and they directly employ models trained on datasets with human-made videos, thereby overlooking the inherent distributional differences between human-made videos and AI-generated videos. Other studies like T2VQAKou et al. (2024) and AIGC-VQA Lu et al. (2024) have implemented end-to-end evaluations of videos, but they only consider video quality and text-video alignment. Therefore, we designed a multi-branch fusion evaluation scheme, which includes a technical quality encoding branch, a video realness encoding branch, and a text-video alignment encoding branch. The technical quality encoder is implemented based on the Swin-3D network Liu et al. (2022b), the video realness encoder utilizes the ConvLexNet-3D network Liu et al. (2022a), and the text-video alignment is achieved using the BLIP network Li et al. (2022b). Considering that technical quality and realness rely solely on visual perception, we prioritize the fusion of quality and realness features before integrating them with text-video alignment features. Additionally, we adopted a progressive training strategy that ensures the independent effectiveness of each branch while promoting collaboration among the branches. Experimental results show that T2VEval effectively assesses the performance of different T2V-generated videos and has broad potential applications, achieved state-of-the-art performance on both T2VEval-Bench and T2VQA-DBKou et al. (2024).
2 RELATED WORK
2.1 Subjective evaluation and dataset for T2V Generation
@@
Chivileva et al. Chivileva et al. (2023) constructed the first T2V generated video evaluation dataset. They obtained 201 textual prompts by combining ChatGPT and manual writing, covering topics such as famous figures, regions, and special festivals. Five T2V models, namely Aphantasia, Text2Video-Zero, T2VSynthesis, Tune-a-Video, and Video Fusion, were selected to generate 1005 videos. Subsequently, 24 subjects were invited to subjectively score the generated videos from the perspectives of naturalness and consistency. Liu et al. proposed FETV Liu et al. (2024b). They designed a set of diverse and fine-grained textual prompts and classified them into three aspects: main content, control attributes, and text complexity. At the same time, a time category was specifically designed for text-to-video, covering spatial and temporal dimensions. On this basis, the authors selected four T2V models, including CogVideo, Text2Video-zero, ModelScopeT2V, and ZeroScope, and conducted subjective evaluations in terms of static quality, temporal quality, overall alignment, and fine-grained alignment. EvalCrafter Liu et al. (2024a) first constructed a comprehensive textual prompt corpus covering daily objects, attributes, and actions. It specified the metadata of common elements in the real world and expanded them using large language models like ChatGPT to obtain complete text prompts. It also enriched and expanded the prompts through those written by real-world users and from text-to-image (T2I) datasets. Subsequently, 2500 videos were generated on 5 T2V models, and subjective evaluations were carried out in terms of video quality, text-video consistency, motion quality, and time consistency. Wu et al. Wu et al. (2024b) constructed the TVGE dataset. They adopted the prompt set proposed in EvalCrafter Liu et al. (2024a) and generated 2543 videos on 5 T2V models. Ten professional scorers were invited to score the videos in terms of video quality and video-text consistency. Huang et al. Huang et al. (2024) believed that the existing evaluation benchmarks could not accurately reflect the highlights and problems of different generation models. Based on this, they proposed Vbench, which further refined the evaluation dimensions of T2V. Sixteen fine-grained evaluation dimensions for generated videos were designed, including subject identity inconsistency, motion smoothness, time flicker, and spatial relationship. One hundred text prompts were designed for each evaluation dimension. In addition, the authors also evaluated T2V models for different content categories, such as people, animals, and landscapes. A total of 8 content categories were considered, and 100 prompts were designed for each category. Based on this, the authors conducted subjective evaluations on 4 T2V models and collected corresponding 16-dimensional scores. Li et al. Zhang et al. (2024c) constructed the LGVQ dataset, which contains 468 text prompts covering combinations of different foreground contents, background contents, and motion states. Six T2V models were used to generate 2808 videos, and 60 subjects were invited to perceptually score the spatial quality, temporal quality, and text-video alignment of each video. Based on the LGVQ dataset, the authors carried out benchmark tests using 14 IQA methods, 16 VQA methods, and 9 methods based on CLIP and visual question answering. Zhang et al. Zhang et al. (2024b) pointed out that the existing subjective evaluation benchmarks for T2V generally have limitations in reliability, reproducibility, and practicability. In view of this, this paper proposes the T2VHE benchmark, which specifies the evaluation dimensions (including video quality, temporal quality, motion quality, text alignment, etc.) used in T2V evaluation, the annotator training process and example descriptions, and a user-friendly GUI interface. The authors further proposed a dynamic scoring scheme. By combining the automatic scoring method to pre-annotate the videos, the amount of manually annotated data was subsequently reduced by 50%. Kou et al. Kou et al. (2024) believed that the current T2V evaluation benchmarks have overly complex divisions of evaluation dimensions and lack a simple and effective evaluation index. Therefore, the authors constructed T2VQA-DB and invited 24 subjects to score 10,000 text-generated videos. The subjects were required to give an overall score considering video quality and text-video consistency.
@@
All of the above works focus on constructing general subjective evaluation benchmarks for text-to-video generation. At the same time, there are also works that conduct subjective evaluations for specific video content types or evaluation angles. Chen et al. Chen et al. (2024b) constructed the GAIA dataset for action quality assessment of generated videos from the perspective of actions in the generated videos, covering various full-body, hand, and facial actions. TC-Bench Feng et al. (2024) focuses on evaluating the temporal composition and consistency of each element in the generated videos. T2VBench Ji et al. (2024) and DEVIL Liao et al. (2024) evaluate from the perspective of the dynamic degree of text-to-video. T2V-CompBench Sun et al. (2024) evaluates T2V models from seven dimensions such as dynamic attribute binding, spatial relationship, and generation computing ability in terms of the ability of T2V models to combine different objects, attributes, and actions. In addition, Miao et al. Miao et al. (2024) paid special attention to the security risks of text-to-video. They pointed out that the generated videos may contain illegal or unethical content, and previous work lacked a quantitative understanding of security. Therefore, the authors introduced T2VSafetyBench to evaluate T2V models in 12 aspects such as violence, gore, discrimination, political sensitivity, copyright, and trademark infringement.
2.2 Objective Metrics for T2V Generation
Chivileva et al. Chivileva et al. (2023) proposed the first T2V evaluation index, which assesses the naturalness of the video and the similarity between the text and the video. In the text similarity evaluation, corresponding captions are generated for each video based on the BLIP model. Then, the similarity between the generated captions and the text prompts is calculated using the text encoder of BLIP. For the evaluation of naturalness, a series of handcrafted features are introduced, including texture, sharpness, color, spectrum, entropy, contrast, keypoint detection, and the number and size of speckles. FETV Liu et al. (2024b) aimed at the problem that previous text-to-video evaluation metrics such as FVD, CLIPSim, and IS could not accurately reflect human subjective perception. It improved FVD and CLIPSim and proposed the index FVD-UMT for evaluating video fidelity and the index UMTScore for evaluating text-video alignment. Wu et al. Wu et al. (2024b) proposed T2VScore to automatically evaluate text-to-video in terms of video quality and text-video consistency. T2VScore consists of two components, T2VScore-A and T2VScore-Q, which evaluate text-video consistency and video quality respectively. Specifically, T2VScore-A uses LLMs to generate question-answer pairs for the prompt and then uses MLLMs to ask questions about the video and answer them. The consistency score is obtained by calculating the answer accuracy. The structure of T2VScore-Q adopts a mixture of experts structure, and the basic architecture adopts Fast-VQA. EvalCrafter Liu et al. (2024a) believes that traditional metrics such as FVD and IS cannot well fit human perception, and simple metrics cannot achieve effective evaluation. The article divides video quality, consistency, motion quality, and time consistency into fine-grained sub-dimensions, obtaining a total of 17 evaluation dimensions, and designs a dedicated evaluation algorithm for each dimension (for example, to evaluate the ability of the current T2V model to generate text, EvalCrafter directly uses the PaddleOCR toolkit Ma et al. (2019) to extract text from the generated video. Then, the word error rate, normalized edit distance, and character error rate are calculated and averaged). VBench Huang et al. (2024) also makes a finer division based on the evaluation of video quality and text-video consistency, finally obtaining 16 evaluation dimensions and corresponding 16 evaluation algorithms. Kou et al. Kou et al. (2024) designed the first end-to-end text-to-video evaluation model T2VQA based on the proposed T2VQA-DB dataset. It includes a video fidelity encoding module based on Siwn-3D and a text-video consistency representation module based on BLIP. Finally, the quality score is regressed through a large language model. Lu et al. Lu et al. (2024) improved on T2VQA by introducing a progressive training strategy, enabling different branches of the model to gradually learn relevant knowledge and achieving SOTA performance on T2VQA-DB. Li et al. Zhang et al. (2024c) found that the current metrics could not comprehensively and effectively evaluate text-to-video based on the LGVQ benchmark test. Therefore, the authors proposed the UGVQ model, which systematically extracts spatial, temporal, and text-video alignment features by combining the ViT model, SlowFast network, and CLIP model, and obtains multi-dimensional scores through MLP regression.
3 T2VEval-Bench: A Multi-Dimensional Benchmark Dataset
3.1 Division of Subjective Evaluation Dimensions
@@
Traditional video quality assessment task Kou et al. (2024); Wang et al. (2019); Sinno and Bovik (2018); Hosu et al. (2017a, b) typically represents video quality using a single score, this approach may obscure the strengths and weaknesses of T2V models in specific tasks. It fails to provide a detailed analysis of the dimensions where the model performs well or poorly. Existing evaluation approaches targeting T2V models also suffer from limited applicability. To address these challenges, we proposed a multi-dimensional subjective evaluation method. Specifically, we decomposed the evaluation of text-to-video generation into four dimensions:
@@
3.2 Compilation and Screening of Prompts
Table 1: Detailed information about the models included in this article, including frame rate, duration, resolution, abilities and release date of the version used in the experiment. *Sora has not yet opened its interface, and only 48 published videos participated in our evaluation.
Model & Version Duration FPS Resolution Abilities
Sora* 8-60s 30fps $1280\times 720$ or $1920\times 1080$ T2V
Runway Gen-3 Alpha 8s 24fps $2816\times 1536$ T2V & I2V & V2V
Kling 1.0 5s 30fps $1280\times 720$ T2V & I2V
Dreamina 1.0 12s 8fps $1680\times 720$ T2V & I2V
TypeMovie 32s 24fps $1024\times 576$ T2V & I2V
Pixeling 5s 24fps $1024\times 576$ T2V & I2V
Dream Machine 1.0 5s 24fps $1360\times 752$ T2V & I2V
Vega AI 4s 12fps $1024\times 576$ T2V & I2V
PixVerse V2 4s 18fps $1408\times 768$ or $1024\times 576$ T2V & I2V
Pika 1.0 3s 24fps $1280\times 720$ T2V & I2V
Show-1 3.63s 8fps $576\times 320$ T2V
VideoCrafter 1.6s 10fps $512\times 320$ T2V & I2V
Open-Sora 1.2 4.25s 24fps $1280\times 720$ T2V & I2V & V2V